OpenStudy (anonymous):

find derivative of y=tan8x + cos1/8x

OpenStudy (anonymous):

derivative of tanx = sec^2 x derivative the cosx = -sinx

OpenStudy (anonymous):

is this \[\tan(8x)+\cos(\frac{1}{8x})\]

OpenStudy (anonymous):

so do i use the product rule

OpenStudy (anonymous):

then you do the chain rule 8sec^(8x)-1/8sin(1/8x)

OpenStudy (anonymous):

if so use the chain rule. the derivative of \[\tan(x)\] is \[\sec^2(x)\] so derivative of \[\tan(8x)=8\sec^2(8x)\]

OpenStudy (anonymous):

8sec^2(8x)-1/8sin(1/8x)

OpenStudy (anonymous):

@anilorap should be should be + not -

OpenStudy (anonymous):

derivative of \[\frac{1}{8x}\] is \[-\frac{1}{8x^2}\]

OpenStudy (anonymous):

Q. is the x in the denominator?

OpenStudy (anonymous):

so first term of your derivative is \[8\sec^2(8x)\] and second term is \[\frac{1}{8x^2}\times \sin(\frac{1}{8x})\]

OpenStudy (anonymous):

oh good point!

OpenStudy (anonymous):

sorry i was assuming it was what i wrote. but if not then i am wrong and anilorap is right!

OpenStudy (anonymous):

X is next to 1/8

OpenStudy (anonymous):

\[\cos(\frac{x}{8})\] or \[\cos(\frac{1}{8x})\]?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then i am wrong and anilorap is correct.

OpenStudy (anonymous):

1\[\div\]8X

OpenStudy (anonymous):

(1/8)x

OpenStudy (anonymous):

forget that last one typed wrong

OpenStudy (anonymous):

well in that case my answer is correct

OpenStudy (anonymous):

ayayay.. is the xx is in the denominator yes,, he is lol

OpenStudy (anonymous):

did u get it?